OpenStudy (anonymous):

Find the exact value of sin1290

OpenStudy (owlcoffee):

\[\sin 1290\]?

OpenStudy (anonymous):

Keep subtracting 360 until the angle is between 0 and 360

OpenStudy (wolf1728):

Or divide by 360 =1,290 / 360 = 3.58333333 1,290 / 360 = 3 plus 210 So, you are to find the sine of 210 degrees

OpenStudy (anonymous):

by using calculator you can find the exact value and also use MATLAB .. but f u want to find it without them! i think that u can divide it by (2*pi) (or 360) and you'll have it..

OpenStudy (minatsukisaya95):

find directly the reference angle from calculator..then turn it into fraction...refer the number ratio at special angle...

OpenStudy (anonymous):

\[\text{Sin}[\text{Mod}[1290,360]{}^{\circ}]=-\frac{1}{2} \]